# Neural Geometry Processing via Spherical Neural Surfaces

### v2

@@ 6 Evaluation
The four analytic shapes, which we designed for this experiment, are: mushroom, flower, bobble, and star. In Figure 6, show results on the mushroom only. Refer to the supplemental, for the analytic parametrizations of all four shapes, and for additional results on the flower, bobble, and star. The advantage of using analytic shapes for evaluation is that we can compute the ground truth for any derived geometric quantity (such as curvature). We used symbolic Matlab to compute the ground truth differential quantities, given the analytic parametrization. We provide text files, in the supplemental, which contain the symbolic expressions.
Evaluation Metrics.
@@
  • NFGP [YBHK21] is a method has similar motivation to ours, in the sense that they fit a neural representation to a mesh and compute differential quantities in order to perform edits in the shape. However, they use a deep-implicit representation, so their differential quantities additionally depend on the representation’s adherance to the Eikonal property, whereas we use an explicit representation.
  • i3D [NSS22] is a deep implicit method also computes differential quantities, and improves the overfitting process itself by encouraging the curvature of the deep implicit to match the computed curvature of the input mesh. Because our method, NFGP, and Monge fitting do not require the curvature information from the mesh, we did not provide the mesh-curvature information to this method either and we found that the method suffered as a result.
@@
As expected, the accuracy of the derived quantities is highest when we fit the SNS directly to the analytic parametrization, and the accuracy of the mesh-fitted SNSs approaches the analytically-fitted SNS as the mesh resolution increases. Since the accuracy of the SNS fitted to the fine-resolution mesh (40962 vertices is consistently close to the accuracy of the analytically-fitted SNS, we argue that fitting the SNS to a reasonably fine mesh is a good way to find a smooth approximation of an underlying smooth surface, when no analytic parametrization is available. We also show the distribution of errors for one of the shapes (mushroom), for SNSs fitted to each resolution of the mesh, and for the three comparison methods: Monge-fitting, NFGP and i3D.
@@
In Figure 10 and 11, we put a test-function (a variable-frequency sine-wave scalar field) on two different surfaces: sphere and Igea, respectively. We fit the SNS to various different meshings of these shapes, and in each case compute the result of the applying the SNS LBO to the scalar field. To fit an SNS to a polygonal mesh, we first convert to a triangle mesh by splitting the faces.
@@
To evaluate the accuracy of our Laplace-Beltrami operator on SNSs, we generated random scalar fields using sine waves, and visually compared the results of applying three different forms of the LBO, on five different surfaces (see Figure 8). To compute $\Delta_{S_{\theta}}f$ using the mean curvature formulation (Equation 16), we computed the gradient and Hessians of the scalar field analytically and combined these with the SNS estimates of the normals and mean curvature. We also computed $\Delta_{S_{\theta}}f$ using the ‘divergence of gradient’ definition (Equation 13) - in this version, all of the differential calculations were carried out by autograd. We show the results using vertex-colors, on a dense icosphere mesh.
Finally, we computed the result of applying the cotan LBO to the scalar field (sampled at the vertices of the dense icosphere mesh). At this resolution, all three versions of the LBO are extremely close, however, we notice that the cotan estimate on the spike ball (column 2) contains a small discontinuity, not present in the SNS estimates. From a smooth shape and a smooth scalar field, the resulting $\Delta_{S}$ should be smooth, which is the case for SNS estimates.
Implementation Details.
All of our networks in our experiments were trained on Nvidia GeForce P8 GPUs. The SNS network is a Residual MLP with input and output of size three, and eight hidden layers with 256 nodes each. The networks used to represent scalar fields (in the eigenfunction and heat flow experiments) are very small Residual MLPs with input size three, output size three and two hidden layers with 10 nodes each. (To make the output value a scalar, we take the mean of the three output values.) For both scalar fields and SNS, we use softplus as the activation function. We use the RMSProp optimizer with a learning rate of $10^{-4}$ and a momentum of 0.9. We trained each SNS for up to 20,000 epochs (eight to ten hours), and fixed the normal regularization coefficient to be $10^{-4}$. The time per epoch increases slightly with the number of vertices in the mesh, but for simpler shapes (such as the sphere) the optimization converges in fewer epochs.
@@
Figure 12: Heat Flow (top) and Mean Curvature Flow (bottom). (Top) We can evolve a given scalar field, $f$, specified over the surface of an SNS, $S_{\theta}$, using the Partial Differential Equation (PDE) $f_{t}=\Delta_{S_{\theta}}f$ (the heat equation). We represent the evolving scalar field as a small MLP, whose weights are ‘finetuned’ for up to 100 epochs after every time step. Darker colors denote low values (cold) with lighter colors being high (hot). (Bottom) Taking advantage of the differentiable nature of our representation, we can also compute a mesh-free Mean Curvature Flow (MCF) in which the Mean Curvature and normals are updated at every iteration (the co-ordinate function evolves according to the PDE $X_{t}=\Delta_{S_{\theta(t)}}X$). We update the Spherical Neural Surface using up to 100 finetuning steps, after every iteration of the flow. This formulation of MCF prevents singularities from forming, without any special handling to prevent them [KSBC12].
Heat Flow and Mean Curvature Flow.
@@ Acknowledgements
The authors wish to thank Johannes Wallner, for his constructive comments on an early version of this work, as well as Uday Kusupati and Karran Pandey, for proofreading. We also thank Morena Protti, for the amazing treefrog mesh. RW was supported by the Engineering and Physical Sciences Research Council (grant number EP/S021566/1).

### v3

@@ 6 Evaluation
The four analytic shapes, which we designed for this experiment, are: mushroom, flower, bobble, and star. In Figure 6, we show results on the mushroom only. Refer to the supplemental, for the analytic parametrizations of all four shapes, and for additional results on the flower, bobble, and star. The advantage of using analytic shapes for evaluation is that we can compute the ground truth for any derived geometric quantity (such as curvature). We used symbolic Matlab to compute the ground truth differential quantities, given the analytic parametrization. We provide text files, in the supplemental, which contain the symbolic expressions.
Evaluation Metrics
@@
  • NFGP [YBHK21] is a method with similar motivation to ours, in the sense that they fit a neural representation to a mesh and compute differential quantities in order to perform edits in the shape. However, they use a deep-implicit representation, so their differential quantities additionally depend on the representation’s adherance to the Eikonal property, whereas we use an explicit representation.
  • i3D [NSS22] is also a deep implicit method that computes differential quantities, and it improves the overfitting process itself by encouraging the curvature of the deep implicit to match the computed curvature of the input mesh. Because our method, NFGP, and Monge fitting do not require the curvature information from the mesh, we did not provide the mesh-curvature information to this method either and we found that the method suffered as a result.
@@
As expected, the accuracy of the derived quantities is highest when we fit the SNS directly to the analytic parametrization, and the accuracy of the mesh-fitted SNSs approaches the analytically-fitted SNS as the mesh resolution increases. Since the accuracy of the SNS fitted to the fine-resolution mesh (40962 vertices) is consistently close to the accuracy of the analytically-fitted SNS, we argue that fitting the SNS to a reasonably fine mesh is a good way to find a smooth approximation of an underlying smooth surface, when no analytic parametrization is available. We also show the distribution of errors for one of the shapes (mushroom), for SNSs fitted to each resolution of the mesh, and for the three comparison methods: Monge-fitting, NFGP and i3D.
@@
In Figure 10 and 11, we put a test-function (a variable-frequency sine-wave scalar field) on two different surfaces: the sphere and Igea, respectively. We fit the SNS to various different meshings of these shapes, and in each case compute the result of the applying the SNS LBO to the scalar field. To fit an SNS to a polygonal mesh, we first convert to a triangle mesh by splitting the faces.
@@
To evaluate the accuracy of our Laplace-Beltrami operator on SNSs, we generated random scalar fields using sine waves, and visually compared the results of applying three different forms of the LBO, on five different surfaces (see Figure 8). To compute $\Delta_{S_{\theta}}f$ using the mean curvature formulation (Equation 16), we computed the gradient and Hessians of the scalar field analytically and combined these with the SNS estimates of the normals and mean curvature. We also computed $\Delta_{S_{\theta}}f$ using the ‘divergence of gradient’ definition (Equation 13) - in this version, all of the differential calculations were carried out by autograd. We show the results using vertex-colours, on a dense icosphere mesh.
Finally, we computed the result of applying the cotan LBO to the scalar field (sampled at the vertices of the dense icosphere mesh). At this resolution, all three versions of the LBO are extremely close, however, we notice that the cotan estimate on the star (row 2) contains a small discontinuity, not present in the SNS estimates. From a smooth shape and a smooth scalar field, the resulting $\Delta_{S}$ should be smooth, which is the case for the SNS estimates.
(vi) What is the effect of different levels of distortion in the spherical mesh embedding?
Figure 12 shows area-distortion colourmaps for four different spherical embeddings of the frog mesh. Spherical embeddings were computed by performing explicit deformations of the mesh-embedding produced by multi-resolution embedding [SPK23]. The distortion colourmap is shown on the original mesh. The area-distortion at a single vertex is computed as the one-ring area on the sphere mesh divided by the corresponding one-ring area on the surface mesh. The second and third rows show the SNS trained on each parametrization, with colouring based on the normals and the mean curvature. We see that first order estimates (normals) are more consistent than second order estimates (curvature). We also see that, although the geometry is close to the input surface in all cases, artefacts such as ‘ripples’ occur more frequently in white areas, where the MLP maps a large area on the sphere to a small area on the surface.
Figure 12: We show area-distortion colourmaps for four different spherical embeddings of the frog mesh, and demonstrate the effect of different patterns of distortion on the accuracy of fitting, and normals and curvature estimates.
Implementation Details
All of our networks in our experiments were trained on Nvidia GeForce P8 GPUs. The SNS network is a Residual MLP with input and output of size three, and eight hidden layers with 256 nodes each. The networks used to represent scalar fields (in the eigenfunction and heat flow experiments) are very small Residual MLPs with input size three, output size three and two hidden layers with 10 nodes each. (To make the output value a scalar, we take the mean of the three output values.) This rather small network-size appears to be sufficient to represent the smallest eigenfunctions, however it might be necessary to add more nodes if we aim to represent highly complex scalar fields. For both scalar fields and SNS, we use softplus as the activation function. We use the RMSProp optimizer with a learning rate of $10^{-4}$ and a momentum of 0.9. We trained each SNS for up to 20,000 epochs (eight to ten hours), and fixed the normal regularization coefficient to be $10^{-4}$. The time per epoch increases slightly with the number of vertices in the mesh, but for simpler shapes (such as the sphere) the optimization converges in fewer epochs.
@@
Figure 13: Heat Flow (top) and Mean Curvature Flow (bottom). (Top) We can evolve a given scalar field, $f$, specified over the surface of an SNS, $S_{\theta}$, using the Partial Differential Equation (PDE) $f_{t}=\Delta_{S_{\theta}}f$ (the heat equation). We represent the evolving scalar field as a small MLP, whose weights are ‘finetuned’ for up to 100 epochs after every time step. Darker colours denote low values (cold) with lighter colours being high (hot). (Bottom) Taking advantage of the differentiable nature of our representation, we can also compute a mesh-free Mean Curvature Flow (MCF) in which the Mean Curvature and normals are updated at every iteration (the coordinate function evolves according to the PDE $X_{t}=\Delta_{S_{\theta(t)}}X$). We update the Spherical Neural Surface using up to 100 finetuning steps, after every iteration of the flow. This formulation of MCF prevents singularities from forming, without any special handling to prevent them [KSBC12].
Heat Flow and Mean Curvature Flow
@@ Acknowledgements
The authors wish to thank Johannes Wallner, for his constructive comments on an early version of this work, as well as Uday Kusupati and Karran Pandey, for proofreading. We thank Morena Protti, for the amazing treefrog mesh. RW was supported by the Engineering and Physical Sciences Research Council (grant number EP/S021566/1). NM was supported by Marie Skłodowska-Curie grant agreement No. 956585, gifts from Adobe, and the UCL AI Centre.